OpenStudy (anonymous):

sec^(2)x -1 can you help prove it? i know its an identity tan^(2)x but can someone help me with the alg

OpenStudy (phi):

did you try writing both sides in terms of sin and cos ?

OpenStudy (anonymous):

1/cos^2x - sin^2x+cos^2x I got there before, but then I dont know where to go from there

OpenStudy (anonymous):

phi help me u know me

OpenStudy (phi):

You have \[ \frac{1}{\cos^2 x} - 1 \] I would try writing 1 as cos^2 x / cos^2 x (so you have a common denominator)

OpenStudy (anonymous):

oh I see 1- cos^2 / cos^2x then 1-cos^2x = sin^2x so sin^2x / cos^2x = tan^2x

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thank you very much :D

OpenStudy (anonymous):

come to my queastion phi please